OpenStudy (anonymous):

Solve negative thirty plus x is greater than or equal to negative fourteen for x. x is greater than or equal to 44 x is greater than or equal to 16 x is less than or equal to 16 x is less than or equal to 44

OpenStudy (ajprincess):

Is this the question? \[-30+x\ge-14x\]

OpenStudy (anonymous):

yes it is

OpenStudy (ajprincess):

(1)\(−30+x\ge−14x\) (2)\(−30+x\ge−14\) Which is ur question (1) or (2)?

OpenStudy (anonymous):

− 30 + x ≥ −14 for x

OpenStudy (ajprincess):

\(-30+x\ge-14\) Add 30 to both sides and see vat u get. can u?

OpenStudy (anonymous):

what does "vat" mean

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

s it x is less than or equal to 16

OpenStudy (ajprincess):

no it is x is greater than or equal to 16. \(-30+x\ge-14\) \(-30+x+30\ge-14+30\) \(x\ge16\)

OpenStudy (anonymous):

oh i see what i did well thanks for ur help

OpenStudy (ajprincess):

yw.:)